OpenStudy (anonymous):

is 2(a-4)=4a-(2a+4) a solution or not a solution

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

what about 5y+2=1/2(10y+4)

OpenStudy (anonymous):

always true

OpenStudy (anonymous):

a solution***

OpenStudy (anonymous):

thank you one more what about 4-d=-(1d-4)

OpenStudy (anonymous):

a solution